OpenStudy (anonymous):

The length of a rectangle is 2x – 3 feet, and the width is 5 feet. What is the area of the rectangle in square feet? A. 7x – 15 B. 10x – 15 C. 10x – 8 D. 7x – 8

OpenStudy (anonymous):

What is the formula for finding area of a rectangle?

OpenStudy (anonymous):

1/2 base time hight

OpenStudy (anonymous):

i dont remember, im practicing for a test, this isnt being turned in. i'm just trying to get the answer so i can try and solve the work on my own.

OpenStudy (anonymous):

that is for triangles...

OpenStudy (anonymous):

yes 1/2(base times height)

OpenStudy (anonymous):

I will tell you the area of a rectangle is L * W you were given L and W Should be easy to figure out..

OpenStudy (anonymous):

oh lol i have dyslexia sorry

OpenStudy (anonymous):

No problem :) Honest mistake

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so whats the correct answer to this question? i want to solve the work on my own for practice.

OpenStudy (anonymous):

no you should be fine :) just delet your answer now. i appreciate your help! ill do the work now for practice

OpenStudy (anonymous):

ok cuz they took my score today and i went from a 60 to a 54

OpenStudy (anonymous):

im trying to bring it bacl up

OpenStudy (anonymous):

wow thats not good, thank you for answering it for me! it looks to be correct!!

OpenStudy (anonymous):

yeah

jimthompson5910 (jim_thompson5910):

As JuanitaM is saying Area = Length * Width A = LW A = (2x-3)*5 A = 5(2x-3) A = ???

OpenStudy (anonymous):

10 x

jimthompson5910 (jim_thompson5910):

that's part of it

jimthompson5910 (jim_thompson5910):

since 5 times 2x = 10x

jimthompson5910 (jim_thompson5910):

there's more to it though

OpenStudy (anonymous):

10x-15

jimthompson5910 (jim_thompson5910):

remember to distribute fully

jimthompson5910 (jim_thompson5910):

much better

OpenStudy (anonymous):

thanks for everyones help! thats what i got too

OpenStudy (anonymous):

ur welcome